What is the pH of 0.23 M HClO2
1
Expert's answer
2020-04-08T13:31:29-0400

Solution:

HClO2 is a weak acid in aqueous media.

Therefore, HClO2 ionizes but less than 100% to produce protons (H+).

The acid dissociation constant for HClO2 is Ka = 1.1*10-2.

Let's set up a reaction table for the following ionization: HClO2 ↔ H+ + ClO2-


Then,

1.1*10-2 * (0.23 - x) = x*x;

x2 + 0.011*x - 0.00253 = 0

Solving this quadratic equation, we get that x = 0.0451

x = [H+] = 0.0451 M.


pH = −log[H+].

Therefore,

pH = −log[H+] = −log[0.0451] = 1.3458 = 1.35

pH = 1.35


Answer: 1.35 is the pH of 0.23 M HClO2.